OpenStudy (anonymous):

@hero Which is the conclusion of this statement? If a and b are both negative then a • b = |a| • |b|. A. If a and b are both negative, a • b = |a| • |b|. B. a • b = |a| • |b| if a and b are both negative. C. a and b are both negative. D. a • b = |a| • |b|. <--- Again two parts, and the conclusion part is D because a and b are both negative. It says it inside the sentence. :) Right? 0_0

hero (hero):

Correct

hero (hero):

It says "If both a and b are negative..." The key word is "If", but a • b = |a| • |b| is in the conclusion portion of the statement.

hero (hero):

So therefore, what you chose for the answer choice is correct.

OpenStudy (anonymous):

@hero Theo finds a large container of batteries in a basement. He randomly selects and tests 14 of the batteries and finds that 4 of them do not work. There are 76 batteries in the container. Which is the best prediction for the number of batteries that will not work? A. 5 B. 18 C. 22 D. 29 76/14=5 5x4=20 22 is the closest, to 20. You'll have to explain it to me if I did it wrong. This one confused me. :)

OpenStudy (anonymous):

So...C?

hero (hero):

Yes C. You can also set up a proportion: \[\frac{4}{14} = \frac{x}{76}\] and solve for x to figure it out.

hero (hero):

Each fraction in the proportion represents \[\frac{\text{no of batteries that don't work}}{\text{total number of batteries tested}}\]

OpenStudy (anonymous):

@hero IDK how to do this one, can you explain? :) Which is a conditional statement for this sentence? In the summertime, we go swimming after lunch. A. If we’re going swimming after lunch, then it is summertime. B. If it is summertime, then we go swimming after lunch. C. If we’re going swimming after lunch, then it is not summertime. D. If it is not summertime, then we don’t go swimming after lunch.

hero (hero):

Which do you believe is correct for this one? In other words, based on the given statement, which choice seems to make the most sense? Process of elimination is a valid reasoning method.

OpenStudy (anonymous):

@hero B sounds the most right. :)

OpenStudy (anonymous):

@hero am i right? :)

OpenStudy (anonymous):

Er..... @Hero ?

OpenStudy (anonymous):

@satellite73 @amistre64

hero (hero):

Oh sorry, yes YOU are correct.

OpenStudy (anonymous):

OMG!! I GOT A 100!! THANK YOU THANK YOU THANK!!

hero (hero):

Congrats